OpenStudy (aravindg):

How to produce a "tick" mark using latex? Also i need to know how to show cancellation of a particular term.

OpenStudy (aravindg):

@ParthKohli ?

OpenStudy (aravindg):

@hartnn

Parth (parthkohli):

\[\checkmark\]

Parth (parthkohli):

\[\not{9}\]or\[\cancel{9}\]I prefer the latter.

Parth (parthkohli):

`\checkmark` to produce the "tick"; `\not{9}` and `\cancel{9}` respectively to show cancellation.

OpenStudy (aravindg):

can you show an example to cancel two linear terms in numerator and denominator of a fraction?

Parth (parthkohli):

Yes, \(\dfrac{\cancel{9}^3}{\cancel{3}^1} = \dfrac{3}{1}\).

OpenStudy (aravindg):

\[\LARGE \checkmark\]

Parth (parthkohli):

`\dfrac{\cancel{9}^3}{\cancel{3}^1}`

OpenStudy (aravindg):

i see.. 1 more query how to put an equation inside a box ?

Parth (parthkohli):

\[\boxed{e^{\pi i}+1 = 0}\] ``` \boxed{e^{\pi i}+1 = 0} ```

OpenStudy (aravindg):

thank you very much :)

Parth (parthkohli):

To conclude everything you have asked till now,\[\cancel{\boxed{\checkmark}}\]:-)

OpenStudy (aravindg):

lols

Parth (parthkohli):

No problem!